OpenStudy (anonymous):

Christy drove 300 miles on her vacation. She drove an average of 1.25 times faster on the second 150 miles of her trip. Which expression represents the time she spent driving? Let x= her speed on the first half of the trip. a.375/x b. 337.5/x c.270x d. 270/x

OpenStudy (welshfella):

on first part of the trip , let time taken = t1 x = 150/t1 on second part, let time taken = t2 1.25x = 150/t2 t1 = 150 /x t2 = 150/1.25x t1 + t2 = 150/x + 150 / 1.25x can you continue from here

OpenStudy (anonymous):

yeah i can continue from there :) thank you soooo much

OpenStudy (welshfella):

yw